# Average Certified Radius is a Poor Metric for Randomized Smoothing

### v1

@@ 5.3 Adaptive Attack on the Sphere
function AdaptiveAdv($f$, $x$, $c$, $\delta$, $T$, $\epsilon$)
$\delta^{*}\leftarrow\delta$
@@
   if $f(x+\delta^{*})\neq c$ then
      break
   end if
   $\delta^{*}\leftarrow$ one step PGD attack on $\delta^{*}$ with step size $\epsilon$
   $\delta^{*}\leftarrow\|\delta\|_{2}\cdot\delta^{*}/\|\delta^{*}\|_{2}$
end for
return $\delta^{*}$
end function
@@
Proof.
$\mathbb{P}(\delta)=\frac{1}{(2\pi)^{d/2}\sigma^{d}}\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i=1}^{d}q_{i}^{2}\right)=\frac{1}{(2\pi)^{d/2}\sigma^{d}}\exp\left(-\frac{1}{2\sigma^{2}}\|\delta\|_{2}^{2}\right).$
This concludes the proof. ∎
(a) Before adaptive attack
(b) After adaptive attack
Figure 5 visualizes the gradient norm distributions for different $p_{A}$ before and after the adaptive attack. We observe that the adaptive attack balances the gradient norm of easy and hard samples. Before the attack, the gradient norm of easy samples is much smaller than that of hard samples, while after the attack, the gradient norm of easy samples is amplified without interfering the gradient norm of hard samples. Therefore, with the adaptive attack, Gaussian training obtains a similar gradient norm distribution to SOTA algorithms, and it can find effective noise samples more efficiently. Pseudocode of the adaptive attack is shown in Algorithm 1, and more detailed description is provided in Section A.2.
5.4 Overall Training Procedure
@@
Input: Train dataset $\mathcal{D}$, noise level $\sigma$, hyperparameters $E_{t}$, $p_{t}$, $m$, $T$, $\epsilon$
Initialize the model $f$
@@
   if epoch $<E_{t}$ then
      Sample $\delta_{1},\dots,\delta_{m}\sim\mathcal{N}(0,\sigma^{2}I)$
      Perform Gaussian training with $\delta_{1},\dots,\delta_{m}$
@@
      if epoch $=E_{t}$ then
         Discard hard data samples in $\mathcal{D}$ with $p_{A}<p_{\rm t}$ to form $\mathcal{D}^{\prime}$
@@
         update dataset weight according to Section 5.2
      end if
      Sample $|\mathcal{D}|$ data samples from $\mathcal{D}^{\prime}$ with replacement to form the train set $\mathcal{D}^{\prime\prime}$
      for input $x$, label $c$ in $\mathcal{D}^{\prime\prime}$ do
         Sample $\delta_{1},\dots,\delta_{m}\sim\mathcal{N}(0,\sigma^{2}I)$
         for i $=1$ to $m$ do
            $\delta_{i}^{*}\leftarrow$ AdaptiveAdv($f$, $x$, $c$, $\delta_{i}$, $T$, $\epsilon$)
         end for
         Perform Gaussian training with $\delta_{1}^{*},\dots,\delta_{m}^{*}$
@@
We now describe how the above three modifications are combined. At the beginning, we train the model with the Gaussian training (Cohen et al., 2019), which samples $m$ noisy points from the isometric Gaussian distribution uniformly at random and uses the average loss of noisy inputs as the training loss. When we reach the pre-defined warm-up epoch $E_{t}$, all data points with $p_{A}<p_{t}$ are discarded, and the distilled dataset is used thereafter, as described in Section 5.1. After this, we apply dataset reweighing and the adaptive attack to training. Specifically, every 10 epoch after $E_{t}$ (including $E_{t}$), we evaluate the model with the procedure described in Section 5.2 and assign the resulting sampling weight to each sample in the train set. In addition, we use the adaptive attack described in Section 5.3 to generate the noisy samples for training. The pseudocode is shown in Algorithm 2.
@@ 6 Experimental Evaluation
We now evaluate our proposed method extensively. Overall, our method always achieves better ACR than SOTA methods, which indicates that focusing on easy data can effectively improve ACR.
Main Result. Table 2 shows the ACR of different methods on CIFAR-10. Detailed description of the training, including applied hyperparameters, is provided in Section A.1. Our method consistently outperforms all baselines on ACR, which confirms the effectiveness of our modification to Gaussian training in increasing ACR. Further, our method successfully increases the certified accuracy at large radius, as we explicitly focus on easy inputs which is implicitly taken by other SOTA methods. Figure 6 further visualizes certified accuracy at different radii, showing that at small radii, SOTA methods (including ours) are systematically worse than Gaussian training, while at large radii, these methods consistently outperform Gaussian training. The success of our simple and intuitive modification to Gaussian training suggests that ACR introduces a systematic bias in method selection, and that the field should re-evaluate RS training strategies with better metrics.
@@
Ablation Study. We present a thorough ablation study in Table 3. When applied alone, all three components of our method improve ACR compared to Gaussian training. Combing two components arbitrarily improves the ACR compared to using only one component, and the best ACR is achieved when all three components are combined. This confirms that each component contributes to the improvement of ACR. In addition, they mostly improve the certified accuracy at large radii and reduce certified accuracy at small radii, which is consistent with our intuition that focusing on easy inputs can improve the ACR. More ablation on the hyperparameters is provided in Appendix B.
7 Discussion
We have shown that ACR does not uniformly represent the robustness of a model for different data. Therefore, the field has to seek better alternative metrics to evaluate the robustness of models under RS. We suggest to use certified accuracy at various radii as a more informative metric, which can be easily computed with the same certification budget as ACR. In addition, since achieving maximum certified accuracy at all radii is a challenging task, we should allow algorithms to customize models for different radii, including the certification hyperparameter $\sigma$. This is similar to the practice in deterministic certified training (Gowal et al., 2018; Mirman et al., 2018; Shi et al., 2021; Müller et al., 2023; Mao et al., 2023; 2024a; 2024b; Palma et al., 2023; Balauca et al., 2024).
While our modifications presented in Section 5 are not designed to improve robustness for the general data distribution, they show effectiveness in increasing certified accuracy at large radius. Other existing algorithms show similar effects. Therefore, it is important to note that the field has indeed made progress over the years. However, new metrics which considers robustness more uniformly should be developed to evaluate RS, and algorithms that outperform generally at various radii are encouraged. We hope this work can inspire future research in this direction.
@@ 8 Conclusion
This work rigorously demonstrates that Average Certified Radius (ACR) is a poor metric for Randomized Smoothing (RS). Theoretically, we prove that ACR of a trivial classifier can be arbitrarily large, and that an improvement on easy inputs contributes much more to ACR than on hard inputs. Empirically, we show that all state-of-the-art (SOTA) strategies reduce the accuracy on hard inputs and only focus on easy inputs to increase ACR. Based on these novel insights, we develop strategies to amplify Gaussian training by reweighing the training data to focus only on easy inputs. Specifically, we discard hard inputs during training, weight the dataset with their contribution to ACR, and apply extreme optimization for easy inputs via adversarial noise selection. With these intuitive modifications to the simple Gaussian training, we replicate the effect of SOTA training algorithms and achieve a new SOTA ACR. Overall, our results suggest the need for evaluating RS training with better metrics.
Appendix A Experiment Details
A.1 Training Details
@@
We use the Certify function in Cohen et al. (2019) to calculate the certified radius, same as previous baselines, where $N=10^{5}$ and $\alpha=0.001$.
@@
A.2 Adversarial Attack Algorithm
@@
where $L$ is the loss function, $f$ is the model, $x$ is the input, $y$ is the label, $\delta$ is the noise, $\delta^{*}$ is the updated noise, sign is the sign function, and $\epsilon$ is the step size. Without specific instructions, our experiments are always conducted with $l_{2}$ PGD. We show the results with $l_{\infty}$ PGD on CIFAR-10 in Section C.2.
Appendix B Additional Ablation Studies
@@
B.1 Effect of Discarding Hard Data
@@
B.2 Effect of Adaptive Attack
@@
Appendix C Additional Results
C.1 Ratio of Remained Data after Discarding Hard Samples
@@
C.2 Experiments on $l_{\infty}$ PGD Attack
@@
The the $l_{\infty}$ attack has the similar effect on results to the $l_{2}$ attack. Increasing the step size or the number of steps of the attack will result in better accuracy at large radii but lower accuracy at small radii.

### v2

@@ 5.3 Adaptive Attack on the Sphere
[…]
@@
[…]
(a) Before adaptive attack
(b) After adaptive attack
@@
The proof of Theorem 2 is deferred to Section A.2. Figure 5 visualizes the gradient norm distributions for different $p_{A}$ before and after the adaptive attack. We observe that the adaptive attack balances the gradient norm of easy and hard samples. Before the attack, the gradient norm of easy samples is much smaller than that of hard samples, while after the attack, the gradient norm of easy samples is amplified without interfering the gradient norm of hard samples. Therefore, with the adaptive attack, Gaussian training obtains a similar gradient norm distribution to SOTA algorithms, and it can find effective noise samples more efficiently. Pseudocode of the adaptive attack is shown in Algorithm 1, and more detailed description is provided in Section C.2.
@@
Input:Train dataset $\mathcal{D}$, noise level $\sigma$, hyperparameters $E_{t}$, $p_{t}$, $m$, $T$, $\epsilon$Initialize the model $f$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
5.4 Overall Training Procedure
We now describe how the above three modifications are combined. At the beginning, we train the model with the Gaussian training (Cohen et al., 2019), which samples $m$ noisy points from the isometric Gaussian distribution uniformly at random and uses the average loss of noisy inputs as the training loss. When we reach the pre-defined warm-up epoch $E_{t}$, all data points with $p_{A}<p_{t}$ are discarded, and the distilled dataset is used thereafter, as described in Section 5.1. After this, we apply dataset reweighting and the adaptive attack to training. Specifically, every 10 epoch after $E_{t}$ (including $E_{t}$), we evaluate the model with the procedure described in Section 5.2 and assign the resulting sampling weight to each sample in the train set. In addition, we use the adaptive attack described in Section 5.3 to generate the noisy samples for training. The pseudocode is shown in Algorithm 2.
@@ 6 Experimental Evaluation
We now evaluate our method proposed in Section 5 extensively. Overall, our method always achieves better ACR than SOTA methods, which indicates that focusing on easy data can effectively improve ACR.
Main Result. Table 2 shows the ACR of different methods on CIFAR-10. Training details, including hyperparameters, are provided in Section C.1. We further include the results of independent runs in Section E.1 to validate the significance of our results. Our method consistently outperforms all baselines on ACR, which confirms the effectiveness of our modification to Gaussian training in increasing ACR. Further, our method successfully increases the certified accuracy at large radius, as we explicitly focus on easy inputs which is implicitly taken by other SOTA methods. Figure 6 further visualizes certified accuracy at different radii, showing that at small radii, SOTA methods (including ours) are systematically worse than Gaussian training, while at large radii, these methods consistently outperform Gaussian training. The success of our simple and intuitive modification to Gaussian training suggests that existing approaches do not fully converge on strategies exploiting the weakness of ACR, and that the field should re-evaluate RS training strategies with better metrics instead of discarding them completely.
Ablation Study. We present a thorough ablation study in Table 3. When applied alone, all three components of our method improve ACR compared to Gaussian training. Combing two components arbitrarily improves the ACR compared to using only one component, and the best ACR is achieved when all three components are combined. This confirms that each component contributes to the improvement of ACR. In addition, they mostly improve the certified accuracy at large radii and reduce certified accuracy at small radii, which is consistent with our intuition that focusing on easy inputs can improve the ACR. More ablation on the hyperparameters is provided in App. D.
@@
7 Beyond the ACR Metric
This work rigorously shows that the weakness of the main evaluation metric ACR has introduced strong biases into the development of randomized smoothing training strategies. Therefore, the field has to seek better alternative metrics to evaluate the robustness of models under RS. When the certification setting (budget $N$ and confidence $\alpha$) is fixed, we suggest using the best certified accuracy at various radii. Specifically, models certified with smaller $\sigma$ tend to achieve higher certified accuracy at smaller radii, whereas models certified with larger $\sigma$ usually have higher certified accuracy at larger radii. To extensively illustrate the power of the developed training strategies and avoid such tradeoff and dependence on $\sigma$, we suggest reporting the best certified accuracy across $\sigma$ at each radius. In other words, this represents the setting where one first fixes an interested radius, and then try to develop the best model with the highest certified radius at the pre-defined radius. Notably, several studies already adopt this strategy (Carlini et al., 2022; Xiao et al., 2022), avoiding the weakness of ACR unintentionally. This evaluation is also consistent to the practice in deterministic certified training (Gowal et al., 2018; Mirman et al., 2018; Shi et al., 2021; Müller et al., 2023; Mao et al., 2023, 2024a, 2024b; Palma et al., 2023; Balauca et al., 2024). Furthermore, the community should also consider randomized smoothing methods with $\sigma$ as a variable rather than a hyperparameter, thus effectively removing the dependence on $\sigma$. While there are some preliminary works in this direction (Nurlanov et al., 2023), they usually break the theoretical rigor, thus more research is needed in this direction.
When the certification setting varies, which is common in applying randomized smoothing to practical tasks, metrics are required to reflect the power of RS training strategies under different certification budgets. To this end, we suggest reporting the empirical distribution of $p_{A}$ for a large $N$, e.g., $N=10^{5}$ as commonly adopted. This approximates the true distribution of $p_{A}$ nicely when $N$ is large, and a higher $p_{A}$ usually leads to a larger certified radius consistently under different certification budgets. To this end, as long as one model has larger $P(p_{A}\geq p)$ for every $p\in[0.5,1]$ than another model, we can claim the former is better consistently in different certification settings. In addition, future research may also report the results under different certification budgets, which can provide insights on the capability of the training strategies under different certification settings as well. With these metrics, practitioners can better understand the capability of RS training strategies under the desired certification setting. A case study is included in App. B to further demonstrate this.
While our modifications presented in Section 5 are not designed to improve robustness for the general data distribution, they show effectiveness in increasing certified accuracy at large radius. Other existing algorithms show similar effects. Therefore, it is important to note that the field has indeed made progress over the years. However, new metrics which consider robustness more uniformly should be developed to evaluate RS, and algorithms that outperform at every radius are encouraged. We hope this work can inspire future research in this direction.
@@ 8 Conclusion
This work rigorously demonstrates that Average Certified Radius (ACR) is a poor metric for Randomized Smoothing (RS). Theoretically, we prove that ACR of a trivial classifier can be arbitrarily large, and that an improvement on easy inputs contributes much more to ACR than on hard inputs. Empirically, we show that all state-of-the-art (SOTA) strategies reduce the accuracy on hard inputs and only focus on easy inputs to increase ACR. Based on these novel insights, we develop strategies to amplify Gaussian training by reweighting the training data to focus only on easy inputs. Specifically, we discard hard inputs during training, weight the dataset with their contribution to ACR, and apply extreme optimization for easy inputs via adversarial noise selection. With these intuitive modifications to the simple Gaussian training, we replicate the effect of SOTA training algorithms and achieve the SOTA ACR. Overall, our results suggest the need for evaluating RS training with better metrics. To this end, we suggest using the best certified accuracy at various radii and reporting the empirical distribution of $p_{A}$ for a large $N$ in replacement of ACR. We hope this work will inspire future research in this direction and motivate the community to permanently abandon ACR in the evaluation of randomized smoothing.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Deferred Proofs
A.1 Proof of Theorem 1
Now we prove Theorem 1, restated here for convenience.
See 1
Proof.
A.2 Proof of Theorem 2
Now we prove Theorem 2, restated here for convenience. See 2
Proof.
$\mathbb{P}(\delta)$ $=\frac{1}{(2\pi)^{d/2}\sigma^{d}}\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i=1}^{d}q_{i}^{2}\right)$
$=\frac{1}{(2\pi)^{d/2}\sigma^{d}}\exp\left(-\frac{1}{2\sigma^{2}}\|\delta\|_{2}^{2}\right).$
This concludes the proof. ∎
Appendix B On the Empirical Distribution of $p_{A}$
In this section, we include a case study to further demonstrate the idea we presented in Section 7 about using the empirical distribution of $p_{A}$ to evaluate the RS model. We first show how to calculate the certified accuracy for every interested radius and certification budget efficiently, given the ECDF of $p_{A}$. Then, we show how to convert an existing certified accuracy-radius curve to the ECDF of $p_{A}$, thus the literature results can be easily reused. These features facilitate practitioners to apply RS models in practice, as the certification budget usually varies in different scenarios.
Figure 7: The empirical distribution of $p_{A}$ on CIFAR-10 with $\sigma=1.0$ and $N=10^{5}$. Note that this curve is independent of $\alpha$.
Figure 8: The certified accuracy-radius curve under $N=100$ and $\alpha=0.01$ on CIFAR-10 with $\sigma=1.0$, converted from Figure 7.
B.1 From ECDF to Certified Accuracy under Different Certification Settings
We consider four RS models trained on CIFAR-10. The ECDF of $p_{A}$ under $N=10^{5}$ is shown in Figure 7. Assume now instead of certifying with $N=10^{5}$ and $\alpha=0.001$, we want to certify with $N=100$ and $\alpha=0.01$, which is less expensive and more practical to run on edge devices. We are interested in the certified accuracy-radius curve under this new certification setting.
Since the ECDF of $p_{A}$ is calculated with a large $N$, we assume it equals the true population of $p_{A}$. When it is computed with a relatively small $N$, the difference between ECDF and the true population might be non-trivial, and corrections should be done in such cases. We leave this as future work, and focus on the case where the ECDF is accurate enough to represent the true population.
We take the case where $r=0.5$ to demonstrate the process. Given $r=0.5$, $N=100$ and $\alpha=0.01$, we run a binary search on $p_{A}$ to identify the minimum $p_{A}$ required to certify the radius. In this case, the minimum $p_{A}$ is $0.81$. Then, we look up the ECDF curve to find the corresponding $P(p_{A}\geq 0.81)$. For our method in Figure 7, $P(p_{A}\geq 0.81)=0.316$. Therefore, the expected certified accuracy under the new certification setting is $0.316$. We note that this is an expected value, as multiple runs can lead to different certified accuracies due to the randomness of the noise samples.
Now, repeat this process for all radii of interest, and we can obtain the certified accuracy-radius curve under the new certification setting. The results are shown in Figure 8. We remark that no inference of the model is conducted in this process, thus the overhead is negligible.
B.2 From Certified Radius to ECDF
Computing ECDF for a given model requires the same computation as computing the certified accuracy-radius curve, as the former is a middle step of the latter. Fortunately, given the certified accuracy-radius curve commonly presented in the literature with large $N$, we can also convert it to ECDF of $p_{A}$.
The idea is to revert the process in Section B.1. Assume we are given the certified accuracy-radius curve when $N=10^{5}$ and $\alpha=0.001$. Given an interested $p_{A}$, we first calculate the maximum radius that can be certified with this $p_{A}$ under this certification setting. Then, we can look up the certified radius for this $p_{A}$ in the curve, and obtain the corresponding certified accuracy. This is the probability of $p_{A}$ being greater than the given $p_{A}$, ignoring the difference between ECDF and true population. Repeat this process for all interested $p_{A}$, and we can obtain the ECDF of $p_{A}$.
Appendix C Experiment Details
C.1 Training Details
@@
We use the Certify function in (Cohen et al., 2019) to calculate the certified radius, same as previous baselines, where $N=10^{5}$ and $\alpha=0.001$.
@@
C.2 Adversarial Attack Algorithm
@@
where $L$ is the loss function, $f$ is the model, $x$ is the input, $y$ is the label, $\delta$ is the noise, $\delta^{*}$ is the updated noise, sign is the sign function, and $\epsilon$ is the step size. Without specific instructions, our experiments are always conducted with $l_{2}$ PGD. We show the results with $l_{\infty}$ PGD on CIFAR-10 in Section E.3.
$\sigma$ ACR 0.00 0.25 0.50 0.75 1.00 1.25 1.50 1.75 2.00 2.25 2.50 0.25 0.5613$\pm$0.0022 76.43$\pm$0.11 68.58$\pm$0.34 58.93$\pm$0.34 48.15$\pm$0.10 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.5 0.7587$\pm$0.0011 59.00$\pm$0.20 54.40$\pm$0.26 49.37$\pm$0.16 44.31$\pm$0.14 39.09$\pm$0.19 34.14$\pm$0.05 28.91$\pm$0.05 22.83$\pm$0.13 0.0 0.0 0.0 1.0 0.8434$\pm$0.0003 41.74$\pm$0.43 39.23$\pm$0.44 36.46$\pm$0.31 33.73$\pm$0.33 30.95$\pm$0.27 28.30$\pm$0.08 25.62$\pm$0.09 23.22$\pm$0.15 20.74$\pm$0.15 18.40$\pm$0.25 16.12$\pm$0.19
Table 5: Mean and Standard deviation of ACR and certified radius on CIFAR-10 using our method. The results are based on three independent runs with different random seeds.
Appendix D Additional Ablation Studies
@@
D.1 Effect of Discarding Hard Data
@@
D.2 Effect of Adaptive Attack
@@
Appendix E Additional Results
E.1 Result Significance
Table 5 shows the mean and standard deviation of the ACR and certified accuracy-radius curve on CIFAR-10 using our best hyperparameters reported in Table 4. The results are based on three independent runs with different random seeds. The standard deviation is relatively small, indicating that the results are stable and reproducible
E.2 Discarded Data Ratio
@@
E.3 Replacing $l_{2}$ with $l_{\infty}$ PGD Attack
@@
The $l_{\infty}$ attack has the similar effect on results to the $l_{2}$ attack. Increasing the step size or the number of steps of the attack will result in better accuracy at large radii but lower accuracy at small radii.